# Quantum Corrections to Symmetron Fifth-Force Profiles

### v1

@@ Acknowledgements
The authors thank Clare Burrage, Ben Elder, Christian Käding and Björn Garbrecht for helpful discussions and comments on this draft. This work was supported by the University of Manchester, the Science and Technology Facilities Council (STFC) [Grant No. ST/X00077X/1], and a United Kingdom Research and Innovation (UKRI) Future Leaders Fellowship [Grant Nos. MR/V021974/1 and MR/V021974/2].
@@ A.1 Eigenvalue Problem
We denote by $\Psi_{\mathbf{n}}(t,\mathbf{x})$ the discrete eigenfunctions of the fluctuation operator L, as defined in (4.2). They satisfy
@@
This particular resolution lends itself to the interpretation that the spatial part of the eigenfunctions of L are simply the eigenfunctions of the “Hamiltonian” $\hat{H}=-\nabla^{2}+U(r)$. Plots of the “central potential” $U$ are given in Figure 14(a) and Figure 14(b).
@@
Figure 14:
Modes with zero or negative eigenvalues are problematic in the Euclidean. The former lead to divergences while the latter indicate instabilities. There is a subtle difference in the nature of the zero/negative mode problem when considered in Minkowski spacetime vs Euclidean space. We have no dispersion relation to constrain $\omega$, so it would appear that the Minkowski space fluctuation operator can only have a continuous spectrum. However, since the spectral representation of the Minkowski space Green’s function has the schematic form
@@
a divergence could still arise if the Euclidean part of the fluctuation operator admits a discrete zero mode. In other words, the problem of zero/negative modes only applies to the Euclidean part of the Minkowski space operator.
In the literature on vacuum decay [52, 62, 40], one tends to show the existence of zero modes by taking the gradient of the equation of motion eq. (3.8). It would appear that we can do the same here by writing
@@
Hereafter, we will work in the thin-wall approximation. Translating to it involves the replacements we made in the earlier section, along with the assignment $r\rightarrow R$ in the centrifugal term [40]. This changes our interpretation of the Schrödinger equation. No longer does it describe a central potential but an effective one-dimensional potential
@@ A.2 General Solution
and is uniquely determined by imposing continuity, continuity of the first derivative and normalisability. The exterior modes satisfy
@@
By changing the independent variable to the normalised external field,
@@
Since $l(l+1)/R^{2}-E_{nl}>-4\gamma^{2}$, $n$ is real. If $n<0$, normalisability requires that we set $B_{l}=0$ for all $l$. If $n$ is positive, we may swap $P_{2}^{n}$ for $P_{2}^{-n}$ (see Section 3) and are still required to set $B_{l}=0$. Furthermore, since $P_{2}^{n}$ for negative $n$ and $P_{2}^{-n}$ for positive $n$ coincide, we may set $n>0$ without loss of generality and represent the exterior part of the eigenfunctions as
@@
Rearranging (A.21) gives the spectrum
@@
We encountered this equation in Section 3 and showed that it is solved by functions resembling hypergeometric functions,
@@
where $z=\chi_{-}^{2}/g^{2}$. Since $F_{-a}$ blows up at the origin of the source, normalisability demands $D_{l}=0$ for all $l$, which leaves
@@ A.3 Continuity, Differentiability and Normalisation
Figure 15: A plot of the $W$ as a function of $n$. The root corresponds to the value of $n$ which satisfies the “quantisation condition”. Note its proximity to $n=1$.
@@
where $W$ denotes the Wrońskian. This is the “quantisation condition” which constrains the allowed values of $n$. It cannot be solved in closed form but is easily approximated numerically. A plot of the Wrońskian at $s=0$ as a function of $n$ is given in Figure 15 for a set of test values of the Lagrangian parameters. We observe one root near $n=1$, omitting the infinite number of roots for negative $n$ which normalisability rules out. Indeed, for a wide variety of physical parameter values, the root is always close to unity. A simple Newton-Raphson iteration is sufficient to arrive at a decent approximation to $n$, and we have plotted the resulting eigenfunction in Figure 16.
In conclusion, the Euclidean part of the fluctuation operator admits one tower of discrete modes with positive eigenvalues. There are no zero/negative modes.
Appendix B Green’s Function for Piecewise-Constant Mass
Here, we give a derivation of the Green’s function for a free field with a piecewise constant mass. The method employed here serves as a basis for the one used in Section 4. Consider the equation
@@ Appendix B Green’s Function for Piecewise-Constant Mass
with $m_{+}<m_{-}$. This equation describes a step-wise free system with an effective mass that depends on the sign of the coordinate. We suppose that the coordinate space Green’s function $G(t,t^{\prime};x,x^{\prime})$ may be written as a weighted superposition of temporal plane waves
@@
and $G^{<}(x,x^{\prime};E)=G^{>}(x^{\prime},x;E)$. The two constituent functions satisfy the homogeneous equation
@@
We shall use the symbol $B$ to denote the bound contribution to the Green’s function. While $0<E<m_{+}$, the spectral sum comprises normalisable eigenfunctions of the differential operator in (B.1). The equation
$\left[-\frac{\text{d}^{2}}{\text{d}x^{2}}+m_{+}^{2}-E^{2}\right]y_{+}(x)=0\>,$ (B.7)
describes both the $x$- and $x^{\prime}$-dependence of $B^{\gtrless}_{l,+}(x,x^{\prime};E)$, the $x$-dependence of $B^{>}_{l,\pm}(x,x^{\prime};E)$ and the $x^{\prime}$-dependence of $B^{<}_{l,\pm}(x,x^{\prime};E)$. The equation
@@
where $k_{-}^{2}=m_{-}^{2}-E^{2}$. The usual homogeneous boundary conditions will require $B^{>}_{l,+}(x,x^{\prime};E)$ to tend to zero as $x\rightarrow\infty$ but leave it unconstrained in $x^{\prime}$, thus
@@
Finally, $B^{>}_{l,-}(x,x^{\prime};E)$ tends to zero as $x^{\prime}\rightarrow$-$\infty$ and is unconstrained in $x$, hence
@@
The Green’s function is continuous in $x$ and $x^{\prime}$, but it’s derivative experiences a jump discontinuity along $x=x^{\prime}$ which is given by
@@
Each pair of bound functions with the same combination of signs in their subscript must satisfy this condition. They may each be solved to give the following parameter values,
@@
Note the resemblance between these parameters and the transmission and reflection amplitudes. The bound Green’s function $B(x,x^{\prime};E)$ is thus completely specified by
@@
While $m_{+}<E<m_{-}$, exterior eigenfunctions look like plane waves, while interior eigenfunctions are still attenuated. Consequently, homogeneous boundary conditions remain for equations defined for $x,x^{\prime}<0$ while the boundary behaviour of solutions defined for $x,x^{\prime}>0$ will, for now, remain unspecified. The equation for $y_{+}(x)$ is now solved by trigonometric functions,
@@
While $m_{-}<E$, both exterior and interior eigenfunctions look like plane waves. Now the equation for $y_{-}(x)$ is also solved by trigonometric functions
@@
where $p_{-}^{2}=E^{2}-m_{-}^{2}$. We adopt the symbol $S$ to denote the scattering contributions to the Green’s function. The form of the solution in the northeast quadrant does not change,
@@
The mixed sign solutions are unconstrained in general, but the falloff condition means the only term which contributes is the outgoing travelling wave,

### v2

@@ Acknowledgements
The authors thank Clare Burrage, Ben Elder, Christian Käding and Björn Garbrecht for helpful discussions and comments on this draft. This work was supported by the University of Manchester, the Science and Technology Facilities Council (STFC) [Grant No. ST/X00077X/1], and a United Kingdom Research and Innovation (UKRI) Future Leaders Fellowship [Grant Nos. MR/V021974/1 and MR/V021974/2]. The Mathematica notebook supporting the results presented in this paper can be found at: DOI:10.5281/zenodo.18431681.
@@ A.1 Eigenvalue problem
We denote by $\Psi_{\mathbf{n}}(t,\mathbf{x})$ the discrete eigenfunctions of the fluctuation operator L, as defined in eq. (4.2). They satisfy
@@
In other words, the spatial part of the eigenfunctions of L are simply the eigenfunctions of the “Hamiltonian” $\hat{H}=-\nabla^{2}+U(r)$. Plots of the “central potential” $U$ are given in figures 13(a) and 13(b).
@@
Figure 13: The effective potential of the classical configuration.
Modes with zero or negative eigenvalues are problematic in the Euclidean version of this calculation. The former lead to divergences while the latter indicate instabilities. There is a subtle difference in the nature of the zero/negative mode problem when considered in Minkowski spacetime vs. Euclidean space. Our dispersion relation does not fix a single value of $\omega$, so it would appear that the Minkowski space fluctuation operator can only have a continuous spectrum. However, since the spectral representation of the Minkowski space Green’s function has the schematic form
@@
a divergence could still arise if the spatial part of the fluctuation operator admits a discrete zero mode. In other words, the problem of zero/negative modes only applies to the spatial part of the Minkowski space operator.
In the literature on vacuum decay [28, 31, 36], one tends to show the existence of zero modes by taking the gradient of the equation of motion (3.8). It would appear that we can do the same here by writing
@@
Hereafter, we will work in the thin-wall approximation. Translating to it involves the replacements that we made in the earlier section, along with the assignment $r\rightarrow R$ in the centripetal term [36]. This changes our interpretation of the Schrödinger equation. No longer does it describe a central potential but an effective one-dimensional potential
@@ A.2 General solution
and it is uniquely determined by imposing continuity, continuity of the first derivative and normalisability. The exterior modes satisfy
@@
By changing the independent variable to the normalised external field
@@
Since $l(l+1)/R^{2}-E_{nl}>-4\gamma^{2}$, $n$ is real. If $n<0$, normalisability requires that we set $B_{l}=0$ for all $l$. If $n$ is positive, we may swap $P_{2}^{n}$ for $P_{2}^{-n}$ (see section B.2) and are still required to set $B_{l}=0$. Furthermore, since $P_{2}^{n}$ for negative $n$ and $P_{2}^{-n}$ for positive $n$ coincide, we may set $n>0$ without loss of generality and represent the exterior part of the eigenfunctions as
@@
Rearranging eq. (A.21) gives the spectrum
@@
We encounter this equation in section B.2 and show that it is solved by functions resembling hypergeometric functions,
@@
where $z=\chi_{-}^{2}/g^{2}$. See eq. (B.57) for the definition of $F_{\pm a}$. Since $F_{-a}$ blows up at the origin of the source, normalisability demands $D_{l}=0$ for all $l$, which leaves
@@ A.3 Continuity, differentiability and normalisation
Figure 14: A plot of the Wrońskian $\left.W\left[F_{a}\left(z(s)\right),P_{2}^{-n}(u(s))\right]\right|_{s=0}$ as a function of $n$. The root corresponds to the value of $n$ which satisfies the “quantisation condition” A.30. Note its proximity to $n=1$.
@@
where $W$ denotes the Wrońskian. This is the “quantisation condition”, which constrains the allowed values of $n$. It cannot be solved in closed form but is easily approximated numerically. A plot of the Wrońskian at $s=0$ as a function of $n$ is given in figure 14 for a set of test values of the Lagrangian parameters. We observe one root near $n=1$, omitting the infinite number of roots for negative $n$ that normalisability rules out. Indeed, for a wide range of physical parameter values, the root is always close to unity. A simple Newton-Raphson iteration is sufficient to arrive at a decent approximation to $n$, and we have plotted the resulting eigenfunction in figure 15.
In conclusion, the spatial part of the fluctuation operator admits one tower of discrete modes with positive eigenvalues. There are no zero/negative modes.
Appendix B Green’s function
This appendix provides a full derivation of the Green’s function quoted at the end of section 4. As a pedagogical introduction to the manipulations involved, we first analyse the simpler problem of a free field with piecewise constant mass in appendix B.1. Then the calculation relevant for the symmetron field is given in detail in appendix B.2.
B.1 Example: piecewise constant mass
First, we consider a 1+1 dimensional system, which could correspond to a field with constant mass in the positive and negative domain but experiences a discontinuous jump at the origin. The Green’s function for such a system satisfies the equation
@@ B.1 Example: piecewise constant mass
with $m_{+}<m_{-}$. This equation describes a step-wise free system with an effective mass that depends on the sign of the coordinate. We suppose that the coordinate-space Green’s function $G(t,t^{\prime};x,x^{\prime})$ may be written as a weighted superposition of temporal plane waves
@@
and $G^{<}(x,x^{\prime};E)=G^{>}(x^{\prime},x;E)$. The two functions satisfy the homogeneous equation
@@
We shall use the symbol $B$ to denote the bound contribution to the Green’s function. While $0<E<m_{+}$, the spectral sum runs over normalisable eigenfunctions of the differential operator in eq. (B.1). The equation
$\left[-\frac{\text{d}^{2}}{\text{d}x^{2}}+m_{+}^{2}-E^{2}\right]y_{+}(x)=0$ (B.7)
describes the $x$- and $x^{\prime}$-dependence of $B^{\gtrless}_{l,+}(x,x^{\prime};E)$, the $x$-dependence of $B^{>}_{l,\pm}(x,x^{\prime};E)$ and the $x^{\prime}$-dependence of $B^{<}_{l,\pm}(x,x^{\prime};E)$. The equation
@@
where $k_{-}^{2}=m_{-}^{2}-E^{2}$. The usual homogeneous boundary conditions will require $B^{>}_{l,+}(x,x^{\prime};E)$ to tend to zero as $x\rightarrow\infty$ but leave it unconstrained in $x^{\prime}$. Thus,
@@
Finally, $B^{>}_{l,-}(x,x^{\prime};E)$ tends to zero as $x^{\prime}\rightarrow$-$\infty$ and is unconstrained in $x$. Hence,
@@
The Green’s function is continuous in $x$ and $x^{\prime}$, but it’s derivative experiences a jump discontinuity along $x=x^{\prime}$, which is given by
@@
Each pair of bound functions with the same combination of signs in their subscript must satisfy this condition. They may each be solved to give the following parameter values:
@@
Note also the resemblance between these parameters and the transmission and reflection amplitudes. The bound Green’s function $B(x,x^{\prime};E)$ is thus completely specified by
@@
While $m_{+}<E<m_{-}$, exterior eigenfunctions look like plane waves, while interior eigenfunctions are still attenuated. Consequently, homogeneous boundary conditions remain for equations defined for $x,x^{\prime}<0$, while the boundary behaviour of solutions defined for $x,x^{\prime}>0$ will, for now, remain unspecified.
The equation for $y_{+}(x)$ is solved by trigonometric functions,
@@
While $m_{-}<E$, both exterior and interior eigenfunctions look like plane waves. The equation for $y_{-}(x)$ is again solved by trigonometric functions
@@
where $p_{-}^{2}=E^{2}-m_{-}^{2}$. We adopt the symbol $S$ to denote the scattering contributions to the Green’s function. The form of the solution in the top-right quadrant of the $x-x^{\prime}$ plane does not change,
@@
The mixed sign solutions are unconstrained in general, but the falloff condition means the only term that contributes is the outgoing travelling wave,